OpenStudy (anonymous):

Simplify by factoring. Assume that all expressions under radicals represent nonnegative numbers. √x^17

OpenStudy (anonymous):

\[\Large \sqrt{x^{17}} =\sqrt{x\cdot x^{16}} =x^8\cdot \sqrt x\]

OpenStudy (anonymous):

because... \[\LARGE \sqrt[m]{x^n}=x^{\frac nm}\] in your case m=2 (but we don't write it...)

OpenStudy (anonymous):

Thanx Kreshnik

OpenStudy (anonymous):

Glad to help :)